### v1

Abstract
Nordic skiing provides fascinating opportunities for mathematical modelling studies that exploit methods and insights from physics, applied mathematics, data analysis, scientific computing and sports science. A typical ski course winds over varied terrain with frequent changes in elevation and direction, and so its geometry is naturally described by a three-dimensional space curve. The skier travels along a course under the influence of various forces, and their dynamics can be described using a nonlinear system of ordinary differential equations (ODEs) that are derived from Newton’s laws of motion. We develop an algorithm for solving the governing equations that combines Hermite spline interpolation, numerical quadrature and a high-order ODE solver. Numerical simulations are compared with measurements of skiers on actual courses to demonstrate the effectiveness of the model.
Keywords.
mathematical modelling, ordinary differential equations, spline interpolation, cross-country skiing, sport science
MSC codes.
65D05, 65L05, 97M10
1 Background
Nordic skiing, also known as cross-country skiing, is a winter sport that attracts participants at all fitness levels, ranging from recreational skiers to athletes competing in World Cup and Olympic races. Nordic ski courses or trails are groomed, snow-covered paths that are carefully designed to provide a variety of terrain interspersing flats with undulating stretches, moderate climbs with steep up/down-hills, and gently winding turns with tight curves – all of which provide both casual skiers and serious athletes with something to enjoy and challenge themselves. Compared to alpine skiing, where the entire course is downhill and gravity does the work of propulsion, Nordic skiers spend a majority of their time on flats and uphills where they must maintain their forward motion by generating propulsive forces using a combination of techniques that engage their entire body, including arms, legs, and torso. As a result, Nordic skiing is frequently touted as an ideal activity to foster aerobic fitness [10, 18] and even reducing mortality [12], with Stöggl al. [23] concluding that “cross-country skiing can be regarded as the gold standard winter time aerobic exercise mode, with a high percentage of muscles in the whole body being activated and the highest VO2max values among all sports.”
Beyond its athletic appeal, Nordic skiing also provides diverse opportunities for mathematical modelling studies that combine aspects from course geometry, biomechanics, skier dynamics (governed by muscle propulsion and snow/air resistance), and race strategy. The complex terrain and highly variable snow conditions on a typical ski course induce skiers to employ a variety of techniques (double poling, diagonal striding, step or skid turns, etc.) that add further complexity [17]. This is in contrast with other sports like running, cycling and swimming, where the technique used varies less within a race and is determined almost solely by race distance rather than variations in the course [11, 13]. In this paper, we present a model grounded in Newtonian physics that captures the balance of forces acting on an athlete skiing in both two and three dimensions, which is formulated as a system of nonlinear ordinary differential equations (ODEs). This problem is suitable for treatment within a variety of undergraduate classes in mathematical modelling, ordinary differential equations, or scientific computing.
Sport scientists have already devoted significant effort to investigating the dynamics of Nordic skiing, leading to many studies in which differential equation models have been used to study questions related to athlete propulsion and optimizing race strategies. Our work is largely based on the model of Moxnes, Sandbakk and Hausken [19] (which we abbreviate as MSH), who derived a system of two ODEs that capture the motion of an athlete along a 2D race course in response to forces of propulsion, gravity, snow resistance and aerodynamic drag. Similar ODE-based models of skier dynamics have also appeared in [4, 5, 17, 20, 24], whereas other researchers have developed much more complicated biomechanical models that capture the detailed arm, leg and body motions employed in specific techniques [2, 7, 18].
Our main goal is to provide a detailed derivation of the coupled ODE system for the 2D MSH model, along with an associated numerical scheme. We propose several modifications and improvements to this model that are easily accessible to students at the undergraduate level:
  • We replace the piecewise linear course geometry from MSH [19] with a smoother Hermite cubic spline interpolant that more accurately captures the shape of a real ski course.
  • We constrain the skier to move along the interpolated path, in contrast with some models that introduce numerical errors appearing as a “drift” away from the actual course.
  • The governing equations are solved using a state-of-the art numerical ODE solver (Matlab’s ode15s) that incorporates adaptive time stepping, error control and event detection.
  • The arc length integral for skier distance is approximated using Simpson’s rule which is a fourth-order numerical quadrature formula.
  • We incorporate a realistic 3D course that includes track curvature, which allows us to simulate different braking techniques used by athletes to safely negotiate tightly-curved downhill sections [3]. This is important in race settings where athletes strive for any possible advantage over their competitors.
Taking advantage of Matlab’s built-in interpolation algorithms and ODE solvers [14], we provide numerical simulations that clearly demonstrate the advantages of these model extensions. Taken together, this study demonstrates many natural connections with standard material taught in the undergraduate math curriculum, in courses such as vector calculus, ordinary differential equations, numerical methods, and mathematical modelling. Consequently, we hope to convince the reader that relatively elementary mathematical concepts and techniques can give rise to novel results that have practical application to real problems in sport.
Exercises, Matlab Code and Data Files
Exercises are distributed throughout this paper to provide opportunities to study missing details, to dig more deeply into examples, and to further explore the material. Solutions to all exercises are provided as Supplementary Material which can be downloaded from this website:
http://www.sfu.ca/~jstockie/skiing/
All data files and Matlab codes used to generate the numerical results are also provided with the Supplementary Materials.
2 Constructing a Smooth Ski Course from Sparse Data
Generating realistic simulations of skier dynamics requires a similarly realistic representation of the path that they ski along. However, real ski course geometries are typically only available in the form of a 2D elevation plot (of height versus distance) or in some cases as a GPS file that defines a relatively sparse sample of waypoints along the course. As a result, before we even start to develop a model, it is important to first grapple with several issues related to representing the course itself:
  • Selecting an appropriate parameterization so that the ski course can be represented (in 2D or 3D) as a continuous real-valued function for use in an ODE model.
  • Generating a suitable smooth approximation for a sparse set of data points that remains as faithful as possible to the actual course underlying the given data.
  • Verifying that the final parameterized curve satisfies the specifications for an official race course, and dealing with any gaps or errors in the data.
2.1 Parameterizing a Ski Course
The path that a skier follows while propelling themselves along a real ski course is a 3D curve such as that depicted in Fig. 1a. This particular plot is generated from GPS data [6] measured on the 4.2 km “Ole” course in Toblach, Italy. This course has been officially certified for competition purposes by FIS (Fédération Internationale de Ski) and has been the location of many high-level races including the 2024 FIS World Cup Tour de Ski. The raw GPS data consists of 58 measurements of latitude $\lambda$, longitude $\psi$, and elevation $z$, which must first be converted to Cartesian coordinates on a flat patch of the Earth’s surface. The curve in Fig. 1a is plotted in terms of $(x,y,z)$, where $z$ is elevation, and $x,y$ lie in the horizontal plane [^4] When latitude $\lambda$ and longitude $\psi$ are measured in degrees, the north-south distance (in m) is $y\approx 111133\psi$ whereas the east-west distance is $x\approx 111413\lambda\cos\left(\pi\lambda/180\right)$. The latter contains an extra cosine adjustment factor because the distance between lines of longitude varies with $\lambda$. These two formulas are actually the leading order terms in an approximation that accounts for the oblate spheroidal shape of the Earth, as explained in the Wikipedia article “Geographic coordinate system.”. For convenience, we also translate coordinates so that the course starts at the origin (highlighted in the plot by a green square). A second view of the course is obtained in Fig. 1b by “unwinding” or straightening this 3D curve and displaying it as a 2D plot of elevation versus a new distance variable $\xi$, which represents the arc length measured along the projected path in the $x,y$-plane (the dotted magenta curve in Fig. 1a). This 2D elevation plot of $(\xi,z)$ is the official method for reporting course elevation profiles in competition documents (private communication, A. Maddox, June 9, 2024).
(a) Ole course (3D view) (b) 2D elevation profile
Refer to caption Refer to caption
(c) Three parameterizations for elevation
Figure 1: Three views of the 4.2 km Ole course in Toblach [6]. (a, Top Left) 3D plot of the skier path defined by 58 GPS points $\vec{r}_{i}$ for $i=0,1,2,\dots,57$ (black points), interpolated by a smooth Hermite spline curve. The start and finish in the left foreground are indicated by a green square and purple circle. The dotted magenta curve is the projection of the course in the horizontal plane. (b, Top Right) An elevation profile plot, which is an “unwound” view of the course in 2D and depicts elevation $z$ as a function of horizontal distance $\xi$, where $\xi$ is arc length along the projected skier path (i.e., the magenta curve from (a)). The dashed grey ovals refer to the zoomed plots in Fig. 2. (c, Bottom) Comparison of three parameterizations for elevation: $\xi$ (red, solid), $\alpha$ (blue, dashed) and $t$ (green, dash-dot), with each parameter normalized to lie in $[0,1]$.
This example highlights the importance of exploiting different parameterizations to represent a skier’s path. In general, any space curve may be expressed in vector form as $\vec{r}(\alpha)=\big{(}x(\alpha),\,y(\alpha),\,z(\alpha)\big{)}$, where $\alpha$ is a parameter that in this paper will be chosen in three ways:
– Time parameterization:
When modelling the dynamics of a skier whose path varies in time $t$, we view the position $\vec{r}(t)=(x(t),y(t),z(t))$ as being parameterized by time.
– Distance parameterization:
Ski course geometry is most commonly available as a 2D plot of elevation versus horizontal distance such as in Fig. 1b, where elevation is viewed as a function $z(\xi)$ parameterized by “projected arc length” $\xi$ (the distance along the skier path projected in the horizontal plane).
– Arbitrary parameterization:
When the course is specified using $N+1$ GPS points with coordinates $(x_{i},y_{i},z_{i})$ for $i=0,1,\dots,N$, it is convenient to base the parameter on the index $i$. In this case, we represent the curve as $\vec{r}(\alpha)=(x(\alpha),y(\alpha),z(\alpha))$ where $\alpha\in[0,1]$ and GPS points correspond to discrete values of the parameter $\alpha_{i}=\frac{i}{N}$ (for simplicity, we use equally-spaced points on the unit interval).
In all three cases, we employ the same notation $\vec{r}=(x,y,z)$ to denote position, and make use of different parameterizations as needed. The effect of different parameter choices is nicely illustrated in Fig. 1c, which displays the projected arc length curve $z(\xi)$ from Fig. 1b along with $z(\alpha)$ and $z(t)$. All three parameters have been re-scaled to lie in $[0,1]$ in order to demonstrate how the “physical” elevation plot $z(\xi)$ becomes distorted when displayed using one of the other two parameterizations.
2.2 Interpolating with Cubic Hermite Splines
A dynamic model for a skier moving along a smooth course should determine the location and speed of the skier at any time $t$. However, a real course geometry is only defined at a discrete set of points, either obtained from a 2D elevation profile plot (like Fig. 1b) or a 3D data file in GPS Exchange Format. Therefore, a fundamental first step in the modelling process is to construct a smooth (2D or 3D) curve that passes through these data points. This is a standard problem in data interpolation.
With a bit of foresight, it should be clear that any model of Nordic skiing will require quantities such as slope (or inclination angle) and curvature at the skier’s position $\vec{r}(\alpha)$, parameterized by $\alpha\in[0,1]$. Slope and curvature depend on the first two derivatives of position, $\vec{r}_{\alpha}$ and $\vec{r}_{\alpha\alpha}$, so that any interpolant we use must be twice differentiable. This suggests using a spline interpolant, which is a piecewise-defined function that interpolates each successive pair of data points with a polynomial. For a 2D course a single interpolant $z=Z(\alpha)$ will suffice, but in 3D we will need to construct three spline interpolants, one for each of the coordinate functions: $x=X(\alpha)$, $y=Y(\alpha)$ and $z=Z(\alpha)$.
The simplest choice of interpolant is a linear spline that connects each pair of points with a straight line segment, which is depicted in Fig. 2 in the zoomed-in views of a short section of the Ole course. This interpolant is continuous but only piecewise differentiable, meaning that the derivative (slope) is only piecewise continuous since it is undefined at the data points! Furthermore, curvature is identically zero along all linear segments, so it is impossible to capture any curvature-dependent effects related to skier turning dynamics.
(a) Elevation view (b) Plan (overhead) view
Refer to caption Refer to caption
Figure 2: Comparison of three splines using both elevation (a,left) and plane views (b,right): linear (dotted black), cubic (dashed blue) and cubic Hermite (solid red). The GPS data points are indicated (with black circles) and each spline is plotted using a fine grid of 4000 equally-spaced $\alpha$-points. Both plots show the same “zoomed-in” section of the Ole ski course, for GPS data points $\vec{r}_{i}$ numbered $i=22,\dots,28$ (this is the circled region in Fig. 1).
Another smoother alternative is the usual cubic spline, which is constructed so that the function and its first two derivatives are continuous at all interpolation points. Fig. 2 displays a zoomed-in view of the cubic spline interpolant for the Ole course data (computed with Matlab’s spline function). The resulting curve is clearly smoother than the linear spline, but reveals a disadvantage of the cubic spline: namely, that fitting a smooth interpolant to oscillatory data like a ski course can generate over- and under-shoots that appear visually inconsistent with the original data. These spurious oscillations arising in polynomial interpolation are referred to as Runge’s phenomenon, or more informally as “polynomial wiggle.” This example highlights a problem that is further exacerbated by the irregular spacing between GPS points, which varies from 13–179 m and is easily seen in Fig 1b. When fitting a cubic spline to such data, the amplitude of these oscillations tend to be magnified where interpolation points are closest. This is most evident at the finish where the points are especially tightly clustered, leading to the unrealistic oscillations shown in the close-up view in Fig. 3.
Figure 3: The final 250 m stretch of the Ole course, emphasizing the spurious oscillations that can be generated by a cubic spline when points are relatively flat and closely-spaced.
Ideally, we seek an interpolant that is smooth (at least differentiable) and also avoids creating any new oscillations. This is precisely the aim of Hermite cubic interpolation (implemented in Matlab’s makima function), which relaxes the requirement that the second derivative be continuous and instead “respects monotonicity and is shape-preserving” [14] [^5] Matlab has two functions that implement Hermite cubic splines: pchip and makima. A detailed comparison of these two interpolants with the usual cubic spline is provided in two Matlab Blog posts by Moler [15, 16]. We choose the “modified Akima” Hermite interpolant makima because (according to Moler) “it produces undulations which find a nice middle ground between ‘spline’ and ‘pchip’,” which in our experience means that it also generates the best results when applied to ski course data.. The Hermite spline interpolant pictured in Figs. 2 and 3 clearly eliminates the over/under-shoots in the cubic spline and more closely follows trends in the data; however, this comes at the expense of losing one degree of smoothness at the data points.
Another side-effect of the highly variable spacing between GPS points is that the equal-$\alpha$ parameterization in Fig. 2a distorts the elevation curve, as shown in Fig. 1c. Consequently, the projected arc length parameterization $z(\xi)$ in Fig. 1b is a more “natural” choice for a ski course because it provides an accurate visual representation of geometric quantities like distance, slope and curvature.
2.3 Computing Distance and Inclination Angle
We are now prepared to address the distinction between modelling racers skiing on 2D and 3D courses: in 2D, where the geometry is given by an elevation profile plot of $z$ versus $\xi$; and in 3D, where $(x,y,z)$ coordinates are obtained from GPS data. In the 2D case, points $(\xi_{i},z_{i})$ can be obtained from an elevation profile plot with a data extraction tool such as WebPlotDigitizer [21], and then used to construct the parametric splines $\Xi(\alpha)$, $Z(\alpha)$. In 3D, the GPS points $(x_{i},y_{i},z_{i})$ are interpolated to obtain the splines $X(\alpha)$, $Y(\alpha)$, $Z(\alpha)$, but then an additional step is required to determine the projected arc length parameter $\xi$. This is easily done using the integral formula
[…]
to compute point values of $\xi_{i}=\xi(\alpha_{i})$, which can then be interpolated with a spline $\Xi(\alpha)$.
Another quantity of primary importance in skiing is the distance along the snow surface, denoted $s$, which can also be expressed as an arc length integral:
[…]
When implementing these distance calculations in Matlab, the integrals in (1) and (2) must be computed numerically. To obtain accurate approximations for $s$ (and also $\xi$ for 3D courses), the coordinate splines ($\Xi,Z$ in 2D, $X,Y,Z$ in 3D) are first evaluated on a fine grid of $\alpha$ points. For the 4.2 km Ole course, we use 840 points which yields an average resolution of 5 m for each $\alpha$-interval. The integrals are then approximated using Simpson’s rule, which is fourth-order accurate and has a particularly simple implementation when the $\alpha$ points are equally spaced (see the Matlab code cumsimpson.m [^7] Conveniently, the spline data structure used in Matlab permits spline derivatives $X_{\alpha}$, $Y_{\alpha}$, $Z_{\alpha}$ appearing in the arc length integrands to be computed exactly using the function fnder.). Simpson’s rule is preferable to Matlab’s built-in trapezoidal rule function trapz because of its much higher accuracy. Once the corresponding fine-grid values of $\xi$ are computed, we can then re-parameterize the curve in terms of $\xi$ by computing a new spline for $Z(\xi)$ (plus $X(\xi)$ and $Y(\xi)$ in 3D). Indeed, the smooth elevation curve plotted in Fig. 1b was generated by first approximating the integral (1) with Simpson’s rule, and then fitting a Hermite spline to the resulting $\xi,z$ data.
One more geometric quantity that is of critical importance for capturing skier dynamics is the angle of inclination $\theta(\xi)$ that the course makes with the horizontal at any location $\xi$ (refer to Fig. 4). This angle can be expressed in terms of $z(\xi)$ by recognizing that it is related to the slope $\frac{dz}{d\xi}$ by
$\theta(\xi)=\arctan\left(\frac{dz}{d\xi}\right).$ (3)
Because we have already built the spline $Z(\xi)$ for elevation, the derivative $Z_{\xi}$ can be computed easily using Matlab’s fnder. Then $\theta$ can be evaluated at spline points and used to construct another spline $\Theta(\xi)$ for the inclination angle. This procedure is implemented in the Matlab code setup2d.m, which constructs the Hermite splines based on 2D elevation data from a CSV file. More details are provided in Section 5 regarding the corresponding code setup3d.m for constructing a 3D parameteric spline representation from GPS data.
In closing, we reiterate that in [19], MSH based the majority of their simulations on a linear spline representation of the course. They compared their results to a single run with a cubic spline interpolant, yielding a finish time roughly 20 s slower, which is a reflection of the fact that a linear spline always yields the shortest possible distance between data points. In their view, the cubic spline result was close but still yielded a poorer match with the measured race time forming the basis for their comparisons. This led them to conclude that their model simulations with the cubic spline were “not more realistic than the chosen piecewise linear track,” despite the fact that any differences could just as easily be attributed to their choice of model parameters. We should also recognize that piecewise linear interpolation of GPS data is inherent in the “homologation” procedure that FIS officials use to measure and certify Nordic race courses (and which is discussed further in the next section). One of our main goals is therefore to investigate how employing a smooth spline interpolant affects the accuracy of model results in comparison with a linear approximation.
Exercise 1.
The model devloped by MSH [19] parameterizes a ski course as $z(s)$ using arc length $s$, and replaces slope $\frac{dz}{d\xi}$ in Eq. (3) with $\frac{dz}{ds}\big{/}\sqrt{1-(\frac{dz}{ds})^{2}}$. Show that these two expressions for slope are equivalent.
Exercise 2.
Interpolate the Ole course data with a quintic (degree 5) polynomial. This can be done using the Matlab function spapi instead of makima. Plot the quintic spline alongside the cubic Hermite interpolant and discuss any “problematic” features of the quintic approximation. Because spapi and makima return different spline data structures, generating plots is easiest with the all-purpose fnplt function.
2.4 Homologation and Data Errors
All ski courses that run high-level Nordic races must adhere to guidelines set out in the FIS Cross-Country Homologation Manual [8]. According to the dictionary, “homologation” is the granting of approval by some official authority. In the context of skiing, homologation refers to the FIS-certified process that any race course must go through before an official competition can be held. The manual sets out constraints on course terrain and stadium layout, including limits on the number, length, and slope of hills that are allowed [^8] An easily-digestible summary of the homologation rules can be found on the Wikipedia page “Cross-country skiing trail”.. These specifications are implemented in the form of an Excel spreadsheet that FIS officials use to perform the necessary calculations [8].
A basic understanding of these course design criteria is useful, especially in light of the reality that any GPS measurements must contain errors, which could cause a spline interpolant to deviate significantly from the certified course. Below is a list of the main homologation criteria, which we number as H1–H6:
  1. H1.
    GPS measurements should be taken every 20 m along a course, or at every point where the gradient changes. Regardless of the GPS resolution, our fine-grid of 5 m for spline calculations was chosen to fall well within this 20 m distance constraint. If we interpret “gradient changes” as points where the slope changes sign, then this is a further justification for using Hermite spline interpolation since it avoids introducing any new points where the gradient might change.
  2. H2.
    The actual length of a race course should not exceed the officially posted distance by more than 10%, nor should it be shorter by more than 5%.
  3. H3.
    The average gradient, including all uphills and downhills, should be in the range of 6–14%. This corresponds to an inclination angle $\theta$ (see Fig. 4) lying between 3.4–8.0, since percentage gradients refer to a fraction of 45 degrees [^9] To convert between percentage gradients and angles, use the fact that the instantaneous slope of a curve $z(\xi)$ is the derivative $\frac{dz}{d\xi}$. Then the percent gradient is $\text{\emph{pct}}=100\frac{dz}{d\xi}$, which gives these two conversion formulas: $\text{\emph{pct}}=100\tan\left(\frac{\pi\theta}{180}\right)$ and $\theta=\frac{180}{\pi}\,\arctan\left(\frac{\text{\emph{pct}}}{100}\right)$. A nice diagram relating $\theta$ and pct can be found on the Wikipedia page “Grade (slope)”..
    (a) Inclination angle (b) Forces on the skier
[…]
    Figure 4: (a) A 2D elevation profile with the course parameterized as $z=z(\xi)$, showing the inclination angle $\theta$, arc length $s$ (or skier distance), and speed $v$. (b) A skier is acted upon by forces due to propulsion ($F_{p}$), gravity ($F_{g}$), drag ($F_{d}$) and snow friction ($F_{s}$).
  4. H4.
    All courses should be roughly balanced between uphills, downhills, and undulating terrain. Hills are classified as having gradients between 9–18%, corresponding to inclination angles of 5.1–10.2, and are further divided by length into short hills (extending for 10–29 m) and long hills (over 30 m). Gradients above 18% are allowed, but only over distances of 10 m or less.
  5. H5.
    The race start and finish should both consist of a straight section roughly 100–150 m long, with the finish having a gentle upward climb with slope 2–4%.
  6. H6.
    To ensure athlete safety, a maximum limit on curvature of downhill sections is imposed in terms of centripetal acceleration, $a_{c}=v^{2}/R$, where $v$ is the skier speed and $R$ is the radius of curvature. A table in [8, p. 26] identifies various limits on $a_{c}$ ranging from 2–25 m/s2 depending on race format.
These design criteria will be discussed further after we have developed an algorithm for approximating the course arc length, slope, and curvature in the following sections. Once we have investigated whether any homologation rules are violated, we will consider whether (and how) the input data should be modified to more accurately capture a realistic race course.
Exercise 3.
Notice in Fig. 1b that the Ole course begins immediately with an uphill gradient, which apparently violates homologation criterion H5. This suggests that perhaps the start area was omitted when GPS measurements were taken. With this in mind, modify setup3d.m to extend the start with a straight, flat section of length 100 m, defined by adding two extra points at the same elevation as the original start. To determine $x,y$ components for these points, use linear extrapolation based on the first two GPS points. Plot the extended course as a 2D elevation profile.
Exercise 4.
Download the GPS file for either the “Altprags Uphill” or “Nathalie” course from the Dolomiti NordicSki website [6]. Verify that at least one pair of GPS points are duplicates (that is, they have identical coordinates $\lambda$, $\psi$, $z$). When this occurs the $\xi_{i}$ are not monotone increasing, which can cause problems! Modify setup3d.m to automatically identify and delete any duplicate points before constructing the spline interpolant.
Exercise 5.
Download the GPS file for one of the FIS-rated courses named “Stephanie” or “Saskia” from the Dolomiti NordicSki website [6]. Run the setup3d.m code with this data and investigate whether the resulting spline satisfies the homologation criteria H1–H5 (don’t worry about curvature and H6 for now).
3 2D Model of Nordic Skiing
3.1 ODE System for Skier Dynamics
We begin by developing a model for skiing along a 2D course where elevation is a function of distance $z(\xi)$, as shown in Fig. 1b. There are several reasons for this choice:
  • The course parameterization is simpler, requiring only a single spline curve, and the governing ODEs are also simpler.
  • Most existing dynamic models of Nordic skiing [5, 17, 19, 24] are also two-dimensional, with the skier assumed to move along the 2D course elevation profile.
  • Starting in 2D avoids the complexity of curvature and braking, allowing us to concentrate on testing and calibrating our results against other models before taking the next step to 3D.
Our aim is therefore to determine the skier position $\vec{r}(t)=\big{(}\xi(t),\,z(t)\big{)}$ as a function of time, with $t$ measured in seconds. The skier speed (in m/s) is given by $\vec{v}(t)=\vec{r}^{\,\prime}(t)=\big{(}\xi^{\prime}(t),\,z^{\prime}(t)\big{)}$, where the “prime” denotes a time derivative, and the scalar speed in the direction of motion is the vector magnitude $v(t)=\|\vec{v}(t)\|=\sqrt{\xi^{\prime}(t)^{2}+z^{\prime}(t)^{2}}$. This requires that speed is always positive ($v>0$), meaning that the skier always moves in the forward direction, which is a reasonable assumption for trained athletes.
The skier is treated as a point mass that moves along the curve defining the course, and is assumed to obey Newton’s second law – “mass times acceleration equals force.” The forces acting on the skier are defined for convenience as forces per unit mass (with units of m/s2) and are assumed to come from four main sources:
  • Propulsion force $F_{p}(v)$: which is generated internally by the skier’s muscles and incorporates various effects such as aerobic and anaerobic metabolism, skiing style (skate or classic), technique (diagonal stride, double pole, offset, one-skate, two-skate, etc.) and possibly extra effects arising from technique or race strategy. We use the following speed-dependent function
    $F_{p}(v)=\frac{P(v)}{mv},$ (4)
    where $m$ is the skier mass (in kg) and $P(v)$ represents the power (in watts or kg m2/s3) that is exerted by the muscles to generate forward motion. The “$\frac{1}{v}$” dependence implies that the skier works hardest to accelerate at low speeds, whereas the power demands on their muscles ease off at higher speeds [^10] Our earlier assumption that $v>0$ ensures that this propulsion force is well-defined. However, we could handle a zero speed by replacing the factor $v$ in the denominator of (4) with $\max(v,\varepsilon)$ for some small cut-off parameter $\varepsilon>0$.. This function was proposed by MSH [19] based on a fit to measurements of Nordic ski racers and the specific form for $P(v)$ will be described later in Section 3.3.
  • Gravitational force $F_{g}(\theta)=-g\sin\theta$: where $g=9.81$ m/s2 is the acceleration due to gravity. Note that $F_{g}$ is positive on downhill sections where $\theta<0$ (acting to speed up the skier), whereas it is negative on uphills where $\theta>0$ (slowing the skier down).
  • Snow friction $F_{s}(\theta)=-\mu g\cos\theta$: where $\mu$ is the (dimensionless) coefficient of dynamic friction. According to [4, 19], typical values of $\mu$ lie in the range $[0.03,0.06]$, although in general it depends on a variety of factors including air/snow temperature, sun exposure, grooming quality, ski wax, etc. Note that friction always opposes motion whether the skier is travelling up/downhill.
  • Aerodynamic drag force $F_{d}(v)=-\beta v^{2}$: which we assume depends quadratically on the speed relative to still air [^11] If there is a significant headwind/tailwind, then $v$ should be replaced with a wind-corrected speed (see Exercise 15).. The parameter $\beta=\frac{\rho C_{d}A}{2m}$ (units of m-1) is proportional to air density $\rho$ (in kg/m3), skier cross-sectional area $A$ (m2), and a dimensionless drag coefficient $C_{d}$. As their speed varies throughout a race, trained skiers will adjust their technique and body orientation so that the area $A$ varies with the skier’s speed. So in practice, we will consider the product $C_{d}A(v)$ to be a given function of speed (discussed in more detail in Section 3.3).
Newton’s second law then requires that the sum of all applied forces must balance the acceleration:
$mv^{\prime}=m\Big{[}F_{p}(v)+F_{g}(\theta)+F_{s}(\theta)+F_{d}(v)\Big{]}.$
It should now be clear why we chose to define forces per unit mass, since this permits us to eliminate the skier mass $m$ from all terms except the propulsion force and obtain the simplified equation
$v^{\prime}$ $=\frac{P(v)}{mv}-g\sin\theta-\mu g\cos\theta-\beta v^{2}.$ (5)
This equation is an ODE for $v$ that depends implicitly on the course geometry through the inclination angle as $\theta=\theta(\xi(t))$, so we still need another equation for the time evolution of $\xi$. The time derivative of $\xi$ is just the horizontal component of the skier velocity (see Fig. 4), which can be written as the ODE
$\xi^{\prime}$ $=v\cos\theta.$ (6)
The inclination angle appearing in the right hand side of both ODEs can be obtained by simply evaluating the spline $\theta=\Theta(\xi)$ constructed in Section 2.3.
This completes the model for a 2D skier, which can be summarized compactly as follows. The skier speed $v(t)$ and projected arc length $\xi(t)$ are integrated in time by solving the two ODEs (5)–(6). Wherever $\theta$ appears in the equations, it is replaced with the corresponding spline $\Theta(\xi(t))$ evaluated at the current value of $\xi$. Furthermore, the elevation $z(t)$ and skier distance $s(t)$ can be determined at any time by evaluating the splines $Z(\xi(t))$ and $S(\xi(t))$ respectively. This is an especially simple and efficient algorithm because all splines are derived from the given course geometry, and so can be precomputed before solving any ODEs. There is an additional advantage that regardless of any numerical errors introduced in the calculation of $v(t)$ and $\xi(t)$, the skier’s position will always lie exactly on the spline used to construct the course.
To obtain a well-posed initial value problem, the above ODEs must be supplemented with initial conditions. Since we shifted the course coordinates so that the start lies at the origin, it is natural to initialize the projected distance to $\xi(0)=0$. For the speed, we take the same initial value $v(0)=2$ used by MSH [^12] A more natural choice of initial condition might be $v(0)=0$. However, we can never allow the skier to be at rest (or moving at very low speed) because the propulsion force has $F_{p}(v)\to\infty$ as $v\to 0$. This is a consequence of MSH fitting their power function $P(v)$ to measurements of athletes who were already in motion. Our choice of $v(0)=2$ m/s is reasonable as long as we recognize that $t=0$ actually corresponds to a few seconds after the race start, once the athlete has gotten up to speed. In practice, taking $v(0)$ much smaller than 2 can cause anomalous behaviour owing to unrealistically large propulsion forces..
Exercise 6.
Verify that the ODE system (5)–(6) is “dimensionally consistent” in the sense that within each equation, all terms have the same physical units.
Exercise 7.
Use dimensional analysis (by applying the Buckingham Pi Theorem) to derive a general form for the propulsion force in Eq. (4), assuming only that $F_{p}$ depends on the quantities ${P}_{\text{\emph{max}}}$, $m$ and $v$.
Exercise 8.
Derive an analytical solution for the ODE (5) on a linear course with constant slope.
  1. (a)
    Find an explicit formula for speed $v(t)$ when the course profile is $z(\xi)=b\xi$ for some constant slope $b$. Assume that the skier applies a constant propulsion force, $F_{p}(v)\equiv F_{po}$, which implies that the MSH power function is linear in the speed, $P(v)\propto v$. In constrast with Eq. (4), this force has no “$\frac{1}{v}$” singularity, so you may also assume that the skier starts from rest with $v(0)=0$.
  2. (b)
    Plot the speed $v(t)$ for an uphill climb with a gradient of $+$10% ($b=0.1$) and assume parameter values $\beta=0.0045$, $g=9.8$, $\mu=0.037$, and $F_{po}=1.4$ from the MSH baseline test (see Table 1 and Fig. 5). Observe that the skier’s speed initially increases but then gradually flattens out over time and approaches a constant terminal value. Estimate this terminal speed and the time required to reach it.
  3. (c)
    Repeat part (b) on a constant downhill gradient of --10%, assuming the skier rests in a tuck position with $\beta=0.0025$ and $F_{po}=0$. Compare your results for the uphill and downhill slopes.
3.2 Augmented ODE System
A closely-related model for skier dynamics in [5, 24] treats $z(t)$ and $s(t)$ as additional solution variables and derives ODEs for them based on the course geometry. The time evolution of $z$ and $s$ is determined solely by the components of velocity in the vertical and tangential directions at any point along the course, leading to these two simple ODEs
$z^{\prime}$ $=v\sin\theta,$ (7)
$s^{\prime}$ $=v,$ (8)
which are supplemented by initial conditions $z(0)=s(0)=0$. Altogether the augmented system (5)–(8) consists of 4 nonlinear ODEs in the unknowns $v(t)$, $\xi(t)$, $z(t)$ and $s(t)$, where the inclination angle is obtained by a spline evaluation $\Theta(\xi(t))$. In contrast with the two-ODE model from the previous section, the elevation $z(t)$ is no longer constrained to lie exactly along the skier path because numerical errors in the solution of (7) cause the skier location to “drift.” Unless these errors are controlled, they can generate significant deviations in the skier location away from the course, with a similar drift occurring in $s(t)$.
3.3 Fitting Parameters to Skier Measurements
Unless indicated otherwise, all of our numerical simulations are performed using “baseline” parameters taken from MSH [19] that are listed in Table 1. In particular, their baseline test subject for both experiments and simulations was a male skier with mass $m=77.5$ kg. The gravitational acceleration is $g=9.81$ m/s2, while the air density is $\rho=1.29$ kg/m3.
The drag coefficient $C_{d}A$ varies throughout a race because athletes adjust their posture between an upright or tuck position that alters their frontal cross-sectional area $A$. At the two extremes, a skier moving at high speed on a steep downhill will fall into a deep tuck position to minimize drag and conserve energy; however, at lower speeds (especially when skiing uphill) the skier must focus more on generating forward motion which requires that they adopt a more upright posture. It then seems reasonable to assume the drag coefficient is a function of speed, so we follow MSH and take $C_{d}A(v)$ to be a piecewise constant function of speed that switches at a given threshold $v\gtrless 10$ m/s between two drag values:
$C_{d}A(v)=\begin{cases}0.55,&v\leqslant 10~{}~{}~{}~{}\text{(upright)},\\ 0.35,&v>10~{}~{}~{}~{}\text{(tuck)}.\end{cases}$ (9)
The final remaining detail is the locomotive power function $P(v)$ introduced in Eq. (4). MSH used extensive measurements of ski racers to propose three different functional forms for $P(v)$, which they compared using numerical simulations. We will adopt their “Model 1” that they used in their baseline simulations
$P(v)={P}_{\text{\emph{max}}}\exp\big{(}\text{--}\,(v/b)^{n}\big{)},$ (10)
which is designed to mimic an athlete who generates the highest power when skiing uphill (at lower speeds) and the lowest on downhills (at high speeds). This reflects the same adjustments in technique for the drag coefficient, where the skier has to use their entire body to work hardest on steep uphill sections and maintain their speed, whereas on steep downhills they shift into a tuck position to conserve energy and minimize drag. An intermediate level of power is expended on undulating terrain, where a relatively high speed can be maintained using less energy-intensive techniques. The parameters in this power equation were estimated by fitting with experimental data to obtain $n=4$, $b=10$, and ${P}_{\text{\emph{max}}}=275$.
Table 1: Parameter values for the baseline test case in MSH [19], reported in SI units.
Parameter Symbol Value or Range Units
Gravitational acceleration $g$ 9.81 m/s2
Air density $\rho$ 1.29 kg/m3
Skier mass $m$ 77.5 kg
Snow friction coefficient $\mu$ 0.037
Drag coefficient $C_{d}A(v)$ Eq. (9) m2
Drag parameter $\beta=\displaystyle\frac{\rho}{2m}\,C_{d}A(v)$ $[0.0029,0.0046]$ m-1
Maximum power ${P}_{\text{\emph{max}}}$ 275 kg m2 s-3
Power parameter $b$ 10 m/s
Power exponent $n$ 4
4 2D Model Simulations
The system of nonlinear ODEs (5)–(6), or its augmented formulation (5)–(8), can be solved easily and accurately using Matlab’s stiff, adaptive ODE solver ode15s. This has been implemented in the code skirun2d.m with the variable neqs set to either 2 or 4 (the number of ODEs) to select the desired formulation. In the following sections, we apply the code to solving three test problems:
  1. 1.
    The MSH baseline simulation [19] on a 4.2 km course using the reduced (2-ODE) model. This provides a convenient test to validate our Matlab code and investigate any differences that arise due to our choice of ODE solver and spline interpolant (linear or cubmic Hermite).
  2. 2.
    Comparing to the experimental data in Welde [25] for skiers on a much longer 15 km course, which provides a further validation for a different course geometry.
  3. 3.
    The 4.2 km Ole course (introduced in Section 2) which is a FIS-certified race course that is used to further “stress-test” the model.
A major advantage of ode15s is that it adaptively adjusts the time step and the order of the numerical method to satisfy user-specified error tolerances (we use AbsTol = RelTol = 1e-8). This ensures that especially in the more difficult sections of the course (where curvature is high, or at spline junctions where derivatives lose smoothness) the accuracy of the numerical solution is maintained, while allowing a much larger time step to be taken in “easier” sections. Furthermore, Matlab’s ODE solvers have built-in event detection that allows the integration to be terminated precisely at the end of the course. This is an advantage over models such as Carlsson et al. [5], who employed a constant-step Runge-Kutta solver that forced them to choose an unnecessarily small time step so that they could terminate their code as close as possible to the finish line.
4.1 Example 1: MSH Baseline Comparison
Here we aim to replicate the results from MSH, who compared experimental measurements with simulations for a single skate skier on an unnamed 4.2 km course. The course is defined by 50 points with $(s,z)$ coordinates taken from a 2D plot of elevation versus skier distance [19, Fig. 1]. The data are easily extracted from the given figure using the online data extraction tool WebPlotDigitizer [21], by first uploading the plot image, then calibrating the coordinate axes, manually selecting data points using mouse-clicks, exporting the $(s_{i},z_{i})$ coordinates to a CSV file, and reading the data into Matlab with csvread.
Note that this is a slightly different scenario than that described in Section 2 for the elevation profile plot, where values of $\xi$ come directly from the plot. Instead, we need to compute the $\xi$ coordinates of each point using a procedure analogous to what was described for the 3D GPS data. To this end, we first construct splines $S(\alpha)$ and $Z(\alpha)$ that interpolate the extracted data points, then use Simpson’s rule to approximate the corresponding $\xi$ values from the integral
[…]
after which the required Hermite splines for $Z(\xi)$ and $S(\xi)$ can be built. The makima interpolant for elevation is depicted in Fig. 5a, alongside the linear spline fit used by MSH.
MSH provide no indication whether their course is FIS-rated, but it is nonetheless interesting to verify whether it satisfies any of the homologation criteria in Section 2.4. Starting with criterion H2 for example, the arc length of the Hermite spline interpolant is 4218 m which sits within 0.5% of the stated 4.2 km length. The average gradient is 8.1% (computed as the mean value of $|\theta|$), which sits within the allowable range of 6–14% (H3). The point-wise gradients also lie between --23 and $+$24%, which exceed the FIS limit of $\pm$18%, but only on short sections and so this is also allowed (H4). Being able to easily test these criteria is a useful by-product of our interpolation approach.
(a) MSH elevation profile (b) Simulated speed
Refer to caption Refer to caption
(c) Force components (d) Zoom on a steep uphill
Refer to caption Refer to caption
Figure 5: Simulations of the MSH baseline case. (a, Top Left) Elevation profile shown in terms of $(\xi,z)$ coordinates (black points) extracted from [19, Fig. 1], and interpolated with a Hermite (red, solid line) and linear spline (black, dotted line). (b, Top Right) Simulated speeds for skiers using the two splines, with the average speed of 5.1 m/s shown as a horizontal line. (c, Bottom Left) Forces (per unit mass) throughout the Hermite spline course, including $F_{p}$, $F_{g}$, $F_{s}$, $F_{d}$ and ${F}_{\text{\emph{tot}}}$. (d, Bottom Right) Zoomed view of the force plot on the steep uphill section with $1000\lesssim\xi\lesssim 1150$, where the skier accelerates to a roughly constant speed after which ${F}_{\text{\emph{tot}}}\approx 0$.
We are now ready to compare our model with the baseline test case from MSH. For a linear spline interpolant, the MSH simulation required a time of 815 s to complete which they found differed from the experimentally measured time by 2 seconds (although they did not state whether the actual time was 813 or 817 s). Applying our Matlab code to solve the ODE system (5)–(6) with the same linear interpolant requires 813 s of skiing time, which is within the range reported by MSH. They also provided no details about their ODE solver, so we can expect that some discrepancies will arise from differences in the numerical approximations being used. One advantage of using a robust ODE solver like Matlab’s ode15s is that it integrates seamlessly through derivative discontinuities in the ODE right hand side arising at spline points.
Although MSH performed most simulations on a piecewise linear course, they did include a single run with a cubic spline interpolant for which their skier takes roughly 20 s longer (or 835 s, estimated from [19, Fig. 16]). In comparison, when we run a second simulation with the Hermite spline interpolant, the skiing time increases to 823 s. This time lies roughly midway between those for MSH’s linear/cubic results, which is to be expected since the cubic spline should give the longest arc length (owing to its more oscillatory nature) and the linear spline the shortest.
This 10–20 s variation in skiing times between the linear and smooth interpolants is a 1–2.5% difference, which seems relatively small and indeed, MSH argued that this difference is negligible. The slight difference in course length can be seen in the elevation plot from Fig. 5a, and the corresponding plot of speed versus time in Fig. 5b shows how $v(t)$ follows the same trend for both courses. There are some small discrepancies as well as a noticeable lag in the makima result by the end of the race, although for all simulations the average speed remained consistent at about 5.1 m/s. An animated view of these results is included as a MPEG video in the Supplementary Materials.
Nevertheless, to obtain realistic simulations in a race context where a second or two can distinguish between first and second place [^13] Consider this quote from Hinder et al. [9]: “During the 2022 Games in Beijing, a 1% improvement in the finishing times of 31 athletes who came in from second to eighth place would have enabled them to win a gold medal instead.”, we should strive for as much accuracy as possible in order to differentiate skier dynamics over a range of course conditions and skiing techniques. Because modifying the spline interpolant requires a trivial one-line change to the code, we argue that it is always preferable to employ a smooth Hermite spline interpolant that provides a more accurate and realistic representation of an actual course.
Finally, we include a plot of the force components during our Hermite spline simulation, along with the total force ${F}_{\text{\emph{tot}}}=F_{p}+F_{g}+F_{s}+F_{d}$. Propulsive and gravitational forces clearly dominate throughout, with the largest peaks in $F_{p}$ arising on steep uphills and coinciding with lower speeds and large (negative) values of $F_{g}$. Conversely, steep downhills lead to much larger $v$, positive $F_{g}$, and lower $F_{p}$. The snow friction force has a smaller magnitude on average, but in contrast with other forces it induces a relatively steady contribution opposing forward motion over the entire course. Aerodynamic drag has the smallest impact on skier dynamics, but does spike over short periods on the steepest downhills. It is interesting to observe how the total force oscillates about zero, and that there are regular time intervals where ${F}_{\text{\emph{tot}}}\approx 0$ corresponding to sections with relatively constant slope (such as the steep uphill section with $t\in[250,310]$ and $\xi\in[1000,1150]$, depicted in Fig. 5d). These are stretches where the athlete is able to work themselves up to a roughly constant speed, after which all forces are in balance and hence acceleration drops to zero.
Exercise 9.
Modify the code skirun2d.m to simulate two idealized hill shapes that approximate what a skier would normally encounter on an actual course.
  1. (a)
    Consider a piecewise linear hill described by $z(\xi)=40+\frac{1}{10}\big{|}\xi-400\big{|}$ for $0\leqslant\xi\leqslant 800$. This is analogous to “gluing together” end-to-end the up/downhill sections from Exercise 8b–c, with each stretch having length $40\sqrt{101}\approx 402$ m. Implementing this in Matlab is easiest using the interp1 function with option ’linear’ to interpolate the 3 data points $(0,0)$, $(400,40)$ and $(800,0)$. Use the same parameters as in Exercise 8 and compare the computed speed and completion time with the exact solution.
  2. (b)
    Next, take a smoothed version of the course from part (a) with $z(\xi)=40\sin\left(\frac{\pi\xi}{800}\right)$, and interpolate using 40 equally-spaced spline points on the interval $0\leqslant\xi\leqslant 800$. Before running any simulations, consider this question: If a skier races on this hill and the one from (a), while applying the same constant propulsion force, which course will they complete faster? Then use the code to simulate a skier on the smooth hill to determine whether your intuition was correct.
Exercise 10.
Investigate how important it is to have a speed-dependent switch in the drag coefficient between tuck and upright positions. Compare the baseline simulation for a skier on the MSH course using $C_{d}A(v)$ from Eq. (9), with a second simulation for a constant $C_{d}A\equiv 0.45$, which is just the average of the tuck/upright values.
Exercise 11.
Consider this quote from Carlsson et al. [4]: “Generally a smaller skier is favored on uphills and the larger skier is favored on dowhills and the flat.” Investigate by varying skier mass $\pm$15 kg from the baseline value and comparing the finish times with those in MSH [19, Fig. 11]. Changing mass by itself ignores the fact that more muscular (heavier) skiers can generate higher peak propulsion force. So, repeat your simulations with ${P}_{\text{\emph{max}}}=275\pm 55$ W, assuming the same two skiers generate $\pm$20of the baseline power. This connection between mass and power is discussed further by Carlsson et al. [5].
4.2 Example 2: Welde’s Measurements of Elite Racers
As a second example, we consider an experimental study by Welde et al. [25] that reports results for 36 Norwegian skiers racing on a 15 km course in Tromsø, Norway, using the classic technique. Their study (which we’ll refer to as the “Welde experiment”) involved skiers with a range of abilities who they classified into two equal-sized groups of “slow” and “fast” skiers. The fast group included several top-10 World Cup finishers, and so this example affords us with an excellent opportunity to further validate our model with high-performance skiers and determine whether the power function (specifically ${P}_{\text{\emph{max}}}$) is able to differentiate elite skiers from other competitors.
The authors supply a 2D elevation plot [25, Fig. 2] for their course, which is divided into three laps of length 5 km each, with laps 1 and 3 being repeated runs of the same loop. The $(\xi,z)$ coordinates can be extracted directly from their course plot using WebPlotDigitizer, in contrast with the $(s,z)$ plot from Example 1. Because this course is so much longer than the previous example, we decided to exploit the “automatic extraction” feature of WebPlotDigitizer rather than manually selecting the points. This generates a CSV file with 203 data points that are used to build the Hermite spline interpolant pictured in Fig. 6a. It is important to note that the Welde elevation plot is under-resolved relative to that from Example 1, not only since the course is over three times longer, but also because the $z$-axis scale in the plot image is so much smaller. The reduced resolution introduces several anomalies in the spline interpolant that are easily identified by comparing laps 1 and 3 in Fig. 6a, and focusing on the sections labelled A,B,C.
(a) Welde elevation profile
Refer to captionLap 1Lap 2Lap 3
(b) Simulated speeds
Figure 6: Simulations of the Welde course, taking all parameters equal to the baseline values except ${P}_{\text{\emph{max}}}=434$, corresponding to the time for the winner of the race. (a, Top) Elevation profile interpolated with the Hermite spline. Numbered circles indicate clear discrepancies between laps 1 and 3, which are identical on the actual course. (b, Bottom) Simulations of the skier’s speed during laps 1 and 3 (with the time for lap 3 shifted so it also starts at $t=0$). The black horizontal lines at $v=6.7$ m/s denote the average lap speed.
Welde’s experiment was based on a high-level race during the Norwegian national championships, but the authors did not indicate whether the course was FIS-certified; nonetheless, we will still discuss whether the homologation requirements are met. Criteria H2 and H3 are easily satisfied since the arc length of the spline interpolant is 15048 m (within 0.3% of the stated length) and the average gradient is 8.4% (similar to MSH). However, the gradient varies between --69 and $+$47% over the course, based on a fine-grid sampling of the spline $\Theta(\xi)$, which indicates some steep up- and downhills that might not pass FIS certification. On closer inspection, none of these gradient anomalies occur at or near spline data points but appear instead as spikes or oscillations near the center of a spline interval. This suggests that these spikes are an artifact of the spline interpolation, which likely arise from data extraction errors, or differentiating the spline, or both. But despite certain fine-scale course features that may not be accurately captured, they are few in number and restricted to relatively small sections, and so we are still confident that the Hermite spline provides a reasonable representation of the actual course.
Table 2: Summary of Welde’s data from 36 racers in a 15 km classic ski race in Tromsø, Norway [25, Supporting data file S1]. The rows list skiers divided into two groups of “slow” and “fast” skiers numbering 18 each, with the winner’s result listed separately. The columns list race time (in s) and average speed (in m/s) for the full course, with speeds for laps 1 and 3 included separately. The final column indicates the power parameter ${P}_{\text{\emph{max}}}$ needed in order that our model simulations match the time and speed to ski the full course.
Skier Full course Lap 1 Lap 3 Model
class time speed speed speed ${P}_{\text{\emph{max}}}$
Slow 2560 5.86 6.22 5.83 346
Fast 2389 6.30 6.62 6.28 392
Winner 2268 6.61 6.95 6.60 434
We now describe some specific race results reported by Welde and summarized in Table 2. They observed that the bottom 18 “slow” skiers completed the race in an average time of 2560 s (with average speed 5.86 m/s) whereas the top 18 “fast” skiers finished in 2381 s (average speed 6.30 m/s). The difference in performance between fast and slow skiers is largely a function of their fitness level, technique used, and more specifically the power levels they can sustain throughout a race. We can incorporate this fast/slow difference in our model by manually adjusting the parameter ${P}_{\text{\emph{max}}}$ in the power function (10) until a simulation yields the desired race time. Using this approach, we find that choosing ${P}_{\text{\emph{max}}}=392$ matches the average time for fast skiers is matched (to within 1 s), whereas the slow time is obtained with ${P}_{\text{\emph{max}}}=346$. Welde also singles out the race winner who finished well ahead of all other competitors with a time of 2268 s, that is captured (to within 1 s) by our model by taking ${P}_{\text{\emph{max}}}=434$. This clearly identifies the winner as an outlier, since the difference in maximum power between the winner and average fast skier is almost as large as the difference between the fast and slow groups. This large discrepancy is most likely attributable to the winner’s exceptional technique, based on Welde’s observation that the winner was the only skier who used double-poling (a very energy-intensive technique) throughout the entire race.
The simulations for the winner are depicted in Fig. 6b in terms of the skier speed during laps 1 and 3, which exhibit noticeable variations especially in the three highlighted sections labelled A,B,C. Since these are identical course sections, these differences are most likely due to variations in the spline interpolants for laps 1 and 3. The average speed is indicated by a horizontal line in Fig. 6b, which is essentially identical at 6.7 m/s for both laps. This should be contrasted with the Welde results in Table 2, where the winner’s average speed drops by a wide margin of 5% between the first and last laps. The discrepancies identified in the speed plots for the two laps are too small and short-lasting to cause such a large reduction in speed. Instead, this slowdown between laps 1 and 3 are described by Welde as “positive pacing” and can be thought of as a consequence of muscle fatigue, in which the athlete slows down throughout a race as they exhaust their energy stores and muscles tire. As currently formulated, our model with constant ${P}_{\text{\emph{max}}}$ is not capable of capturing this fatigue effect. But we could emulate racer fatigue in a naïve way by allowing the peak power to decrease with time (or energy expenditure) throughout the simulation. A simple approach to incorporating this effect is explored in Exercise 12, but detailed experimental investigations of racer fatigue would be required in order to build a realistic model that convincingly captures Welde’s results.
Exercise 12.
Notice from Table 2 that skiers are 5–6% slower on lap 3 compared to lap 1, which suggests they may be experiencing fatigue owing to their muscles tiring throughout the course. A simple way to incorporate this effect is to assume that halfway through the race (for $\xi\geqslant 7500$ m), maximum power drops from ${P}_{\text{\emph{max}}}$ to $c{P}_{\text{\emph{max}}}$, where the constant $10<c<$ 1 can be thought of as a “fatigue factor”. Modify skirun2d.m to incorporate muscle fatigue in this manner. Then adjust the values of $c$ and ${P}_{\text{\emph{max}}}$ until the total race time for the “winner” is the same as in Table 2, and their speed on lap 3 is 5–6% slower than on lap 1.
4.3 Example 3: Simulating the Toblach Ole Course in 2D
One of the strengths of our modelling approach is that it can easily handle courses specified as 2D elevation plots or 3D GPS data. Back in Section 2, we introduced the 4.2 km Ole course in Toblach, imported from a GPS file [^14] Toblach (or Dobbiaco in Italian) is located in the South Tyrol region of the Italian Alps (or Dolomites), and lies within a UNESCO World Heritage Site containing 18 major peaks, that also contains the winter resort town of Cortina d’Ampezzo. Cortina hosted the Winter Olympic Games in 1956, and will do so again for the second time in 2026.. This is one of several FIS-rated courses for which a file of GPS waypoints is posted on the Dolomiti NordicSki web site [6]. We explained how to convert GPS data $(\lambda,\psi,z)$ to Cartesian coordinates, which can then be approximated by spline interpolants $X(\xi)$, $Y(\xi)$, $Z(\xi)$ that are parameterized by projected distance $\xi$. The Ole course provides an ideal opportunity to test our 2D model on a course derived from real three-dimensional data, and then after extending the model to 3D later in Section 5, we can compare with simulations that take into account inherently 3D effects such as braking on steep downhills.
After computing the arc length and inclination angle (using the code setup3d.m), it is easy to verify that the Ole course satisfies the homologation criteria H2–H4. The length of the spline is 4157 m and the average gradient is 7.9%, both of which are well within the stated limits (for H2 and H3). The pointwise gradients also remain between $\pm$18as required by H4, except for a few short sections where it spikes to 40%. Regarding criterion H5, the last 100 m of the course are almost completely flat which is not ideal but may still be considered acceptable; however, the start begins immediately with a 5 uphill slope that clearly violates H5 (this issue was addressed previously in Exercise 3).
Simulating the Ole course using the baseline parameters leads to a skiing time of 819.9 s corresponding to an average speed of 5.07 m/s, which is very close to what we computed on the 2D MSH course in Example 1. This is not surprising because, for two courses based on similar design criteria, the skier’s speed should be largely a function of power output, and ${P}_{\text{\emph{max}}}$ is the same in both examples. Fig. 7 depicts the elevation, speed and projected arc length as functions of time $t$. This view of the solution makes it easy to connect steep up/downhill sections with intervals where speed is lower/higher. Also, the progress of skier distance $\xi$ is fairly consistent and doesn’t deviate very far from the straight line with slope equal to the average speed, but the most rapid changes in $\xi$ are easily connected with steep downhills (and slowest changes with uphills).
Figure 7: Simulation of the 4.2 km Ole course with the baseline parameters, displaying speed and projected arc length (top plot, blue/black lines) and elevation (bottom plot, red line). The grey dotted line in the top plot is $\xi(t)=5.07\,t$, corresponding to a skier moving at the constant average speed.
Exercise 13.
On the Dolomiti NordicSki website [6], find the elevation profile plot for the 1.4 km “Stadium Track” course (also called “Albert Sprint”), which is a sequence of two hills having roughly similar size and shape. Use WebPlotDigitizer [21] to extract a sample of $(\xi,z)$ points with roughly 100 m spacing, write them to a CSV file, and use them to simulate a race with the baseline parameters. Next, consider a second manufactured course built from a sum of two “Gaussian bump” functions of the form $a\exp\left(-\frac{1}{c}(\xi-b)^{2}\right)$, and tune the parameters $a,b,c$ to most closely match the first course. Generate CSV data with similar resolution and repeat the simulation. How do the two results compare?
Exercise 14.
Carlsson et al. [5] developed a model very similar to ours, and used it to simulate skiing over a simple hill that consists of a moderate uphill slope (with constant inclination angle $\theta=+3^{\circ}$) followed by a steep downhill (with constant $\theta=-20^{\circ}$), and a smooth transition between the two. Extract data from the elevation plot in [5, Fig. 3], making sure that the gradient at both start/finish is zero. Then simulate a skier using these parameters: skier mass $m=80$ kg, snow friction $\mu=0.03$, and maximum power ${P}_{\text{\emph{max}}}=366.5$ W. Plot the solution and show that the skier completes the course in roughly 136 s. Then compare with the time required to run the course in reverse (right-to-left), starting instead on the steep uphill section. Is this consistent with your intuition?
Exercise 15.
Investigate skiing in windy conditions, and assume there is a constant head/tailwind with speed ${v}_{\text{\emph{wind}}}$. Modify the aerodynamic drag force appropriately, and then repeat the simulations with the ${v}_{\text{\emph{wind}}}$ values displayed in MSH’s Figure 14. When skiing against a headwind, at which value of ${v}_{\text{\emph{wind}}}$ does the aerodynamic drag force start to dominate over other forces?
5 Skiing in Three Dimensions
5.1 Curvature and Braking Force
We now extend the 2D model to a more realistic setting that captures the dynamics of an athlete skiing along a 3D course, who is turning from left to right on top of dealing with changes in elevation on hilly sections. We have already explained in Section 2 the algorithms behind constructing the Hermite spline approximation of a 3D course. Recall that such a course can be written as a vector function $\vec{r}(\alpha)=\big{(}x(\alpha),\,y(\alpha),\,z(\alpha)\big{)}$, where we will employ two choices for the parameter: $\alpha=\xi$ when referring to the course geometry, and $\alpha=t$ when formulating the ODEs governing skier dynamics. The diagram in Fig. 8 depicts an idealized course $\vec{r}(\xi)$ along with its projected path in the $x,y$-plane, and highlights an arbitrary point where the skier is moving with speed $v$ in the tangent direction. The inclination angle $\theta$ is measured relative to the horizontal as in 2D, but must be viewed within the vertical plane containing the tangent vector. In 3D, we require an additional azimuth angle $\varphi$ that captures the horizontal skiing direction measured relative to the $x$-axis. Note that $\varphi$ has the same definition as in 3D spherical coordinates, whereas the inclination angle is related to the usual spherical polar angle $\widetilde{\theta}$ by $\theta=\frac{\pi}{2}-\widetilde{\theta}$. Using standard formulas from vector calculus, the angles may be expressed in terms of the coordinate functions as [^15] A detailed discussion of spherical angles can be found in the Wikipedia article “Spherical coordinate system”.
$\theta=\arctan\Bigg{(}\raisebox{5.69046pt}{$\displaystyle\frac{z_{\xi}}{\sqrt{x_{\xi}^{2}+y_{\xi}^{2}}}$}\Bigg{)}\quad\text{and}\quad\varphi=\operatorname{sign}(y_{\xi})\arccos\Bigg{(}\raisebox{5.69046pt}{$\displaystyle\frac{x_{\xi}}{\sqrt{x_{\xi}^{2}+y_{\xi}^{2}}}$}\Bigg{)}.$ (11)
When building our spline approximations in 3D, we proceed as in 2D and use a fine $\xi_{i}$ grid to compute values for both $\theta_{i}$ and $\varphi_{i}$, and then construct the corresponding Hermite interpolants $\Theta(\xi)$ and $\Phi(\xi)$. These splines can be precomputed and then cheaply evaluated to obtain the two direction angles during a simulation using the current value of $\xi(t)$.
Figure 8: Definition of the inclination angle $\theta$ (measured relative to the horizontal plane) and azimuth angle $\varphi$ (in the $x,y$-plane, measured relative to the $x$-axis) for a 3D ski course parameterized as $\vec{r}(\xi)=\big{(}x(\xi),\,y(\xi),\,z(\xi)\big{)}$. The radius of curvature $R=\frac{1}{\kappa}$ is also shown as the radius of the circle (dashed, green) that best approximates the curve at any given point.
A major difference between skiing in 2D and 3D is that the athlete has the freedom to turn left or right in 3D, thereby introducing an extra centripetal force $ma_{c}$ that acts radially outward (and perpendicular to the tangent direction). This force is proportional to centripetal acceleration $a_{c}={v^{2}}/{R}$, where $R$ is the radius of curvature at any location along the course; because $R$ is related to curvature $\kappa$ by $R=\frac{1}{\kappa}$, an alternate formula is $a_{c}=\kappa v^{2}$. To counteract the centripetal force and avoid sliding sideways, the skier must apply an equal and opposite force to their skis directed radially inward. In all but the most curvy downhill sections of the course, this force induced by $a_{c}$ is very small and easily countered through minor (mostly unconscious) adjustments in technique. However, on the steepest and tightest curves (with large $v$ and small $R$), centripetal effects can become large enough that the athlete can no longer prevent slipping and risks losing control. How a skier handles such turns can be of critical importance in a race as described by Sandbakk et al. [22]: “During the 15-km pursuit race in the 2010 Olympics in Vancouver, Marit Bjørgen passed Justina Kowalczyk in the last downhill turn to win the gold medal.”
In steep downhill turns, skiers reduce their speed by applying a braking force using one of two primary techniques called the skid turn and step turn [^16] In a skid turn, skis are held parallel at a slight angle to the direction of motion, and ski edges are scraped against the snow to both decelerate and turn. In a step turn, a skating motion is combined with rapid side-to-side stepping to maintain speed as much as possible during the turn and dissipate less energy. There is a third technique called the snowplow, which is an extreme version of the skid turn and is more of a beginner’s strategy that is seldom applied by experienced racers. An in-depth discussion of turning techniques can be found in the Nordic Ski Lab article “Nordic downhill ski techniques for safety and speed”.. In both cases, the braking turn generates a force opposing forward motion, which we assume following Sandbakk et al. [22] is proportional to centripetal acceleration so that $F_{b}=\gamma a_{c}=\gamma\kappa v^{2}$, where $\gamma$ is a dimensionless braking coefficient.
It is unrealistic for an athlete to apply a braking force throughout an entire race, even if the force is mostly very small. Therefore, $F_{b}$ should be “thresholded” so that it is zero except within those downhill sections where centripetal acceleration exceeds some braking threshold ${a}_{\text{\emph{c,min}}}$:
$F_{b}=\begin{cases}\gamma\kappa v^{2},&\text{if $\theta<0$ and $\kappa v^{2}>{a}_{\text{\emph{c,min}}}$,}\\ 0,&\text{otherwise.}\end{cases}$ (12)
The specific choice for parameters $\gamma$ and ${a}_{\text{\emph{c,min}}}$ will be discussed separately in Section 5.2.
Now, the only missing detail is a formula for curvature which is a standard identity from vector calculus: [^17] An alternate formula that parameterizes curvature as a function of time is obtained by replacing $\xi$ with $t$ in Eq. (13) and substituting $\vec{r}^{\,\prime}(t)=\vec{v}(t)$, to get $\kappa(t)=\|\vec{v}\times\vec{v}^{\,\prime}\|\big{/}v^{3}$. However, it is Eq. (13) that we implement in our numerical simulations.
$\kappa(\xi)=\frac{\displaystyle\left\|\vec{r}_{\xi}\times\vec{r}_{\xi\xi}\right\|}{\displaystyle\left\|\vec{r}_{\xi}\right\|^{3}}.$ (13)
The derivatives $\vec{r}_{\xi}$ and $\vec{r}_{\xi\xi}$ appearing in (11) and (13) are easily computed as derivatives of the Hermite splines $X(\xi)$, $Y(\xi)$ and $Z(\xi)$ using Matlab’s fnder function. A Hermite spline interpolant is then built for $\kappa=K(\xi)$, similar to what was done for the elevation and azimuth angles.
We can now generalize the 2D model equations (5)–(6) to 3D by adding the braking force to the ODE for $v$ and multiplying the $\xi$ equation by the corresponding angular factor from the spherical coordinate transformation (remembering also to replace $\sin\widetilde{\theta}=\sin(\frac{\pi}{2}-\theta)=\cos\theta$). Then the 3D model equations are
$v^{\prime}$ $=\frac{P(v)}{mv}-g\sin\theta-\mu g\cos\theta-\beta v^{2}-F_{b}(v),$ (14)
$\xi^{\prime}$ $=v\cos\theta,$ (15)
where spline evaluations are used to find $\theta$, $x$, $y$, $z$ and $s$ at the current value of $\xi$. Similar to 2D, we can formulate an augmented version of this system that evolves position variables with four additional ODEs:
$x^{\prime}$ $=v\cos\theta\cos\varphi,$ (16)
$y^{\prime}$ $=v\cos\theta\sin\varphi,$ (17)
$z^{\prime}$ $=v\sin\theta,$ (18)
$s^{\prime}$ $=v,$ (19)
where the azimuth angle is obtained by evaluating the spline $\Phi(\xi)$. In practice, we prefer simulations based on the simpler two-ODE model (14)–(15) because it is more efficient and also constrains the skier to move exactly along the parameterized curve describing the course. Both the two- and six-ODE formulations are implemented in the code skirun3d.m, and although we do not advocate using the augmented model, we still include it for completeness since it has been implemented by others such as [20].
Before moving on, we should mention that some Nordic ski models [5, 20, 24] incorporate an extra contribution to snow friction arising from the component of the centripetal force acting within the $\xi,z$-plane along the direction of travel [^18] Vector calculus tell us that the (signed) curvature for a 2D space curve $z(\xi)$ is given by $\kappa_{{}_{2D}}(\xi)=z^{\prime\prime}(\xi)/(1+z^{\prime}(\xi))^{3/2}$. Note that this formula is equivalent to Eq. (13), except that it is a signed quantity that is positive/negative depending on whether the 2D ’ curve is concave up/down., which is always directed normal to the snow surface. MSH observed that they “find the greatest differences between the simulation and the experimental data on terrain with rapidly changing curvature” [19], which we also see at isolated instants in time. However, because most ski courses have a balance between up/down-hill sections, the +/+/-+ / - contributions from this curvature term tend to cancel out on average, leading to a negligible net effect on the overall skiing time. Therefore, we have chosen to ignore the effects of this vertical curvature on snow friction (which is also consistent with the MSH model [19]).
5.2 Estimating the Braking Parameters
The specification of the braking force in Eq. (12) is incomplete without estimates for the braking coefficient $\gamma$ and threshold ${a}_{\text{\emph{c,min}}}$. The coefficient $\gamma$ can be determined using experimental data reported in Sandbakk et al. [22], who studied elite female skiers applying both skid and step turn techniques while navigating curved downhill sections of a ski course. These authors measured the change in mechanical energy ($\Delta e$) between the start/end of a turn, average skier speed ($\overline{v}$), and “trajectory” or distance travelled ($D$). Their data are summarized in the first four rows of Table 3 for four different combinations of turn geometry ($R$, $D$) and skid/step technique.
Table 3: Measured data from [22] on elite female skiers (with average mass $m=60$ kg) performing several different skid/step turns. The first 4 rows list the original data while the last row contains our estimate of the braking coefficient $\gamma$ based on Eq. (20). Skiers employed the skid technique in the two tighter turns ($R=9,12$), and the step technique in the wider turns ($R=12,15$).
Param. Units Skid turns Step turns
$R=\displaystyle\frac{1}{\kappa}$ m 9 12 12 15
$D$ m 16.5 21 21 24
$\overline{v}$ m/s 6.32 6.61 7.63 7.99
$\Delta e$ kg  m2/s2 26.2 25.4 14.9 17.2
$\gamma$ 0.0060 0.0055 0.0024 0.0028
$\underbrace{\text{\hskip 28.45274pt}}$${\text{\hskip 28.45274pt}}$
${\gamma}_{\text{\emph{skid}}}\approx 0.0058$ ${\gamma}_{\text{\emph{step}}}\approx 0.0026$
Assuming that the braking force is constant throughout each turn, the mechanical energy (or work) can be estimated using the basic physical principle “work equals force times distance”:
$\underbrace{~{}\Delta e{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\big{|}}}_{\text{work}}\;=\;\underbrace{\big{(}m\gamma\kappa\overline{v}^{2}\big{)}}_{\text{force}}\;\cdot\underbrace{~{}D{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\big{|}}}_{\text{distance}}\qquad\Longrightarrow\qquad\gamma=\frac{R\,\Delta e}{m\overline{v}^{2}D}\raisebox{2.84544pt}{.}$ (20)
Substituting the average skier mass $m=60$ kg and other parameters from Table 3 into Eq. (20) yields $\gamma$ values reported in the final row of the table. Turning technique clearly has a significant effect on the braking coefficient, with $\gamma$ for the step turn equal to roughly half that for the skid turn. On the other hand, $\gamma$ depends only very weakly on the turn radius $R$ and length $D$, so we choose the average of the two braking coefficient values ${\gamma}_{\text{\emph{skid}}}=0.0058$ and ${\gamma}_{\text{\emph{step}}}=0.0026$ for the skid- and step-turn techniques, respectively.
The other parameter in our braking model is the threshold ${a}_{\text{\emph{c,min}}}$, which cannot be determined from [22] since they did not consider the different situations when braking turn techniques could be applied. However, the Homologation Manual provides a list of maximum limits for centripetal acceleration [8, p. 26] that are permitted within the various Nordic ski race formats. The smallest two limits allowed in all races are $a_{c}=2$ and 5 m/s2, which suggests two possible values for ${a}_{\text{\emph{c,min}}}$ that will be investigated in the next section.
6 Example 4: 3D Simulations of the Ole Course
We are now ready to simulate a skier moving along an actual 3D ski course. We return to Example 3 and the 4.2 km Ole course from Toblach, and solve the ODE system (14)–(15) by making use of all Hermite spline computed by setup3d (not just those for $Z$, $S$ and $\Theta$). Six simulations are performed taking values of the braking coefficient $\gamma=0$ (no braking), $0.0026$ (step turn), and $0.0058$ (skid turn), each being applied with two threshold values ${a}_{\text{\emph{c,min}}}=2,5$. Otherwise we use the same MSH baseline parameters from Table 1 that were used in 2D simulations.
The corresponding completion times are summarized in the first two rows of Table 4 (labelled “Ole”) where it is clear that using either the skid or step turn increases the finish time by a small margin, on the order of one second or less. The slowest time results from applying the skid turn with threshold ${a}_{\text{\emph{c,min}}}=2$ (requiring an extra 1.12 s) whereas the skier is slightly faster when applying a larger threshold of ${a}_{\text{\emph{c,min}}}=5$. Note that the total time required without braking matches the 2D result of 819.9 s reported in Section 4.3, because the 2D and 3D ODE systems are identical when $\gamma=0$. For comparison purposes, we repeat these same six simulations using GPS data from another 3.9 km FIS-rated course from Toblach called “Stephanie,” which features more tight curves than the Ole course. The corresponding times are given in the last two rows of Table 4, which are also only slightly slower than the “no brake” case; however, the differences are slightly larger than the Ole course since there are more tight curves on which braking forces are applied.
Table 4: Simulated race times for the Ole and Stephanie courses in Toblach. The results are presented for three choices of braking technique (no braking, step turn, skid turn) and two braking thresholds (${a}_{\text{\emph{c,min}}}=2,5$). The figures in parentheses indicate the corresponding changes relative to the time without braking.
Course ${a}_{\text{\emph{c,min}}}$ No brake Step turn Skid turn
Ole (4.2 km) 2 819.94 820.29 ($+$0.35) 820.71 ($+$0.77)
5 820.14 ($+$0.2) 820.39 ($+$0.45)
Stephanie (3.9 km) 2 801.57 802.60 ($+$1.03) 803.92 ($+$2.35)
5 802.45 ($+$0.88) 803.58 ($+$2.01)
Returning to the Ole results, Fig. 9 provides a more detailed look at the effects of braking by plotting the difference in projected distance, $\xi_{2D}-\xi_{3D}$, between the case with no braking ($\xi_{2D}$) and that with skid/step turns ($\xi_{3D}$). The tightest downhill curves occuring in the final third of the course can be identified with the largest discrepancies between the three results. At the end of the course, the skier using skid turns is 7 m behind the idealized athlete who applies no braking, whereas using step turns leaves the skier only 3 m behind. This emphasizes how a simple adjustment in braking technique on only a few difficult turns could allow one skier to make a significant gain over another. As mentioned earlier in Section 5.1 (in reference to Bjørgen’s 2010 gold medal performance), such small differences can be of critical importance in a race scenario where competitors are skiing neck-and-neck and must jockey for any possible advantage.
Refer to captionABC
Figure 9: The difference in projected arc length $\xi$ between the 2D and 3D simulations on the Ole course with three choices of braking parameter ($\gamma=0,{\gamma}_{\text{\emph{step}}},{\gamma}_{\text{\emph{skid}}}$). The braking threshold is set to ${a}_{\text{\emph{c,min}}}=2$ and all other parameters are taken from the MSH baseline test.
Next, we focus on a single simulation with parameters $\gamma=\gamma_{skid}$ and ${a}_{\text{\emph{c,min}}}=2$. The resulting braking force is plotted on top of the elevation profile in Fig. 10a, which underscores how the tightest curves are concentrated within the second half of the course. Plots of the time variation of four solution variables ($F_{b}$, $v$, $\kappa$ and $\theta$) are also given in Fig. 10b, which illustrate that a combination of large speed and curvature is required to generate a significant braking force. In the following discussion, we focus especially on the three shaded regions of the course in Fig. 9 that are labelled A,B,C, and likewise highlighted in Fig. 10b:
(a) Braking force along the 3D course
Refer to caption (b) Solution components versus time (c) Zoomed curvature in Region B Refer to captionABC Refer to caption
Figure 10: (a, Top) The braking force is displayed along the scaled course profile in a 3D view of the solution with $\gamma={\gamma}_{\text{\emph{skid}}}$ and ${a}_{\text{\emph{c,min}}}=$2. The skier brakes much more often during the second half, where the course is mostly downhill. (b, Bottom Left) Plots of $F_{b}$, $v$, $\kappa$ and $\theta$ versus time demonstrate how the braking force arises from a combination of skier speed and track curvature. (c, Bottom Right) Zoomed-in view of the curvature plot in region B.
  • Region A: This starts with a tight curve on an uphill climb at $t\approx 50$ s, followed by a steep downhill that also ends on a highly curved section. Here, the skier only brakes near the bottom of the hill because they are not travelling fast enough on the initial uphill to exceed the braking threshold.
  • Region B: This stretch contains the steepest downhill which also has two tight curves near the hill crest and base, coinciding with the two spikes in $\kappa$ near $t\approx 572$ and $595$ s. The first curve occurs just over the crest when the skier is only beginning to speed up but curvature is large enough to exceed the braking threshold and generate a very small spike in the braking force. A similarly sized “blip” in $F_{b}$ occurs at $t\approx 590$, after the course has flattened but the skier is still travelling at very high speed. This is still before the reaching the tight curve at $t\approx 595$, but there is an extended section beforehand with small-to-moderate curvature that is sufficient to also pass the braking threshold (see Fig. 10c). This example demonstrates how large speed (or curvature) on its own can generate a small braking force.
  • Region C: This is a curvy and mostly downhill portion of the course, which generates the biggest collection of spikes in $F_{b}$ where the braking threshold is exceeded. As a result, we also see a relatively large discrepancy in the distance in Fig. 9 in this section.
An MPEG video of the 3D course is included in the Supplementary Materials.
By adding curvature into our model, we have shown how to quantify the effects of applying various braking techniques. When looking only at the effects on the total race time, the braking force may seem almost negligible on courses like Ole, but our study shows that choosing a more effective turning technique in a few strategic locations can still provide a gain on the order of a few meters. Additionally on more convoluted courses like Stephanie, this impact could be even greater. This simple addition to the model provides opportunities for collaboration with coaches, athletes, and sports scientists to test our model predictions against real athlete data, and to investigate different race strategies with a view to optimizing race results.
Exercise 16.
The emphasis so far has been overwhelmingly on expert skiers, so let’s consider recreational or beginner Nordic skiers, by assuming these changes to the MSH baseline parameters:
  • Beginners ski at much lower power levels, with ${P}_{\text{\emph{max}}}$ being less than half the baseline value.
  • They always remain upright, never switching to a tuck position on downhills.
  • They apply the “snowplow” technique on all downhills regardless of speed or curvature, which has a much larger braking parameter of roughly $\gamma=10{\gamma}_{\text{\emph{skid}}}$.
Adjust the parameters for a skier on the Ole course, and aim for a finish time that is twice the baseline result.
Exercise 17.
Download GPS data for the 1.4 km “Albert Sprint” course (also called “Stadium Track”) from the Dolomiti NordicSki website [6] and simulate a skier in 3D with curvature and braking. Compare your results to the 2D simulation from Exercise 13.
7 Discussion
We have developed an ODE-based model that captures the essential dynamics of Nordic skiing, and provides a stimulating example of the mathematical modelling process that can be easily incorporated into undergraduate mathematics classes. This model is an excellent illustration of the inherent complexity of real problems, which invariably require a multi-pronged modelling strategy that combines concepts and techniques from several different areas including:
  • Data analysis and cleaning: to investigate and correct the random noise and other measurement errors that are inherent in real input data.
  • Function approximation: using piecewise spline interpolation to construct smooth representations for sparse data.
  • Ordinary differential equations: that are derived by combining physical intuition with basic laws of Newtonian mechanics.
  • Numerical algorithms: to approximate the governing equations and validate the model predictions.
  • Scientific visualization: although this was not explicitly addressed, we deliberately employed a varied selection of graphical representations for 2D and 3D data, with the intent to accentuate any differences and to provide extra insight into the results.
Besides the educational value of this skiing model in a classroom setting, we have also attempted to demonstrate how even fairly elementary mathematical techniques can stimulate novel advances in sports science. One example is our use of Hermite splines to approximate a ski course, which provides an attractive balance between providing smoothness and avoiding spurious oscillations. There are also very few 3D models of Nordic skiing, and ours is the first to incorporate the inherently 3D effect of braking in tight downhill curves, which are important strategic sections of competitive races. Examples like these can be exploited to motivate students that the mathematics they are learning has practical value in solving real problems, but can also lead to advances in fundamental knowledge.
Another significant contribution of this paper is our open-source Matlab code, which is flexible enough to simulate skiers on both 2D and 3D course geometries, and handles multiple input formats for the course data. Because Matlab is so widely used in university settings, access to such a code should allow instructors to easily incorporate this material into their courses, while also providing students with a realistic and stimulating problem to experiment on and gain insight from. Besides its educational uses, we also anticipate that this skier simulation code will be of benefit to sports science researchers, who are free to experiment with and generalize the underlying algorithms for their own purposes.
Finally, we should emphasize that this work is still only a preliminary step towards developing a truly realistic model for a very complex and multi-faceted sport, and there are a host of opportunities for further work. One specific area is the study of athlete fatigue in medium- to long-distance events, and how to incorporate fatigue into the power model (beyond the naïve approach we suggested in Example 12). Another natural extension is to extend our simplistic approach for handling braking turns to incorporate “look-ahead,” where a skier anticipates arriving at a tight curve by initiating their braking technique well in advance. There are many other fascinating aspects of Nordic skiing that we have so far ignored, such as the metabolic processes behind muscle power generation, or the variability in course conditions due to weather-induced snow conditions, or track grooming, or athlete race strategy, to name just a few. In terms of strategic decision-making, mathematics is especially well-suited to drive advances in understanding how specific choices of technique or training regimen could optimize race performance under a variety of parameter regimes. Existing work on applying methods from optimization to maximize performance in running races [1, 11] could be extended naturally to Nordic skiing along similar 4lines to what was already done in [7].
Acknowledgements
We thank Kayden Sim (GIS Specialist, Forsite Consultants) and Jake Weaver (Head Coach, Hollyburn Cross-Country Ski Club, North Vancouver, BC). We also appreciate the advice of FIS Technical Delegates and Course Designers John Aalberg (Sooke, BC) and Al Maddox (Atikokan, ON), who assisted us with details regarding course homologation.
Supplementary Materials
  • S1
    Exercise solutions (suppl1Solutions.pdf, 10.3 Mb)
  • S2
    Summary of Matlab codes and data files (suppl2CodesAndData.pdf, 140 Kb)
  • S3
    Matlab codes (suppl3Matlab.zip, 13 Kb)
  • S4
    Data files (suppl4DataFiles.zip, 12 Kb)
  • S5
    Movie of 2D MSH course simulation (suppl5Movie2dMSH.mp4, 2.1 Mb)
  • S6
    Movie of 3D Ole course simulation (suppl6Movie3dOle.mp4, 2.7 Mb)